OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20

OpenStudy (anonymous):

@divu.mkr is it 9 and -15?

OpenStudy (anonymous):

9

OpenStudy (jdoe0001):

\(\bf |2x + 6| - 4 = 20 \implies |2x + 6| = 20+4 \implies \begin{cases} +1(2x + 6) = 20+4\\\quad \\ \bf -1(2x + 6) = 20+4 \end{cases}\)

OpenStudy (jdoe0001):

same gig as before

OpenStudy (anonymous):

2x+6-4=20 2x+2=20 x+1=10 hence x=9

OpenStudy (anonymous):

i know it's 9 what is the second one?

OpenStudy (anonymous):

yep you are correct

OpenStudy (anonymous):

replace in place of + as - for -2x-6-4=20 -2x-10=20 -x-5-10=0 -x-15=0 hence x=-15